# True Positive Rate FPR False Positive Rate

### v1

@@ Abstract
Efficient multi-messenger observations of gravitational-wave candidates from compact binary coalescence (CBC) candidate events relies on data products reported in low-latency by the International Gravitational-wave Network (IGWN). While data products such as HasNS, the probability of at least one neutron star, and HasRemnant, the probability of remnant matter forming after merger, exist, these are not direct observables for a potential kilonova. Here, we present new kilonova light curve and ejecta mass data products derived from merger quantities measured in low latency, by marginalizing over our uncertainty in our understanding of the neutron star equation of state (EoS) and using measurements of the source properties of the merger, including masses and spins. Two additional types of data products are proposed. The first is the probability of a candidate event having mass ejecta ($m_{\mathrm{ej}}$) greater than $10^{-3}M_{\odot}$, which we denote as HasEjecta. The second are $m_{\mathrm{ej}}$ estimates and accompanying ugrizy and HJK kilonova light curves predictions produced from a surrogate model trained on a grid of kilonova light curves from POSSIS, a time-dependent, three-dimensional Monte Carlo radiative transfer code. We are developing this data products in the context of the IGWN low-latency alert infrastructure, and will be advocating for their use and release for future detections.
@@ 1 Introduction
’s (Aasi et al, 2015), ’s (Acernese et al, 2015) and KAGRA’s (Akutsu et al., 2021) fourth observing run (O4) is underway as of May 23, 2023111https://observing.docs.ligo.org/plan, and the search for gravitational wave events and their counterparts (Abbott et al., 2020a) has resumed. Searches for kilonovae are challenging due to the fact that they are short lived events, can be relatively faint, and may not be well localized. Sky localizations for gravitational wave events can span $\approx 100-10,000\ \mathrm{deg^{2}}$ (Röver et al., 2007; Fairhurst, 2009, 2011; Grover et al., 2014; Wen & Chen, 2010; Sidery et al., 2014; Singer et al., 2014; Berry et al., 2015; Essick et al., 2015; Cornish & Littenberg, 2015; Klimenko et al., 2016). However, in spite of these challenges, it is imperative to locate the transient as quickly as possible in the hope of observing the peak of emissions. A number of wide-field survey telescopes are used to try and cover these large sky localizations, such as: the Panoramic Survey Telescope and Rapid Response System (Pan-STARRS) (Morgan et al., 2012), Asteroid Terrestrial-impact Last Alert System (ATLAS) (Tonry et al., 2018), the Zwicky Transient Facility (ZTF) (Bellm et al., 2018; Graham et al., 2019; Ahumada et al., 2024), and in the near future BlackGEM (Bloemen et al., 2015), the Large Synoptic Survey Telescope (LSST) (Ivezic2014), the Nancy Grace Roman Space Telescope [^2] https://roman.gsfc.nasa.gov/ (Andreoni et al., 2024), and Ultraviolet Explorer (UVEX) (Kulkarni et al., 2021). While GW178017 is the only CBC merger event that has provided us with joint observations of a GW signal, kilonova, and short gamma-ray burst (sGRB), it is also potentially possible to identify kilonovae associated with sGRBs (Tanvir et al., 2013; Ascenzi et al., 2019; Jin et al., 2020; Rastinejad et al., 2022), or even serendipitously in survey operations (Andreoni et al., 2021; Almualla et al., 2021).
@@ 2.4 Dynamical and disk wind ejecta from fits: BNS and NSBH
$\dfrac{M^{dyn}_{\rm ej}}{10^{-3}M_{\odot}}=\mathrm{max}\left(0,\left[\dfrac{a}{C_{1}}+b+\left(\dfrac{m_{1}}{m_{2}}\right)^{n}+cC_{1}\right]+[1\leftrightarrow 2]\right)$ (5)

### v2

@@ Abstract
Efficient multi-messenger observations of gravitational-wave candidates from compact binary coalescence (CBC) candidate events rely on data products reported in low-latency by the International Gravitational-wave Network (IGWN). While data products such as HasNS, the probability of at least one neutron star, and HasRemnant, the probability of remnant matter forming after merger, exist, these are not direct observables for a potential kilonova. Here, we present new kilonova light curve and ejecta mass data products derived from merger quantities measured in low latency, by marginalizing over our uncertainty in our understanding of the neutron star equation of state (EoS) and using measurements of the source properties of the merger, including masses and spins. Two additional types of data products are proposed. The first is the probability of a candidate event having mass ejecta ($m_{\mathrm{ej}}$) greater than $10^{-3}M_{\odot}$, which we denote as HasEjecta. The second are $m_{\mathrm{ej}}$ estimates and accompanying ugrizy and HJK kilonova light curves predictions produced from a surrogate model trained on a grid of kilonova light curves from POSSIS, a time-dependent, three-dimensional Monte Carlo radiative transfer code. We are developing these data products in the context of the IGWN low-latency alert infrastructure, and will be advocating for their use and release for future detections.
@@ 1 Introduction
’s (Aasi et al, 2015), ’s (Acernese et al, 2015) and KAGRA’s (Akutsu et al., 2021) fourth observing run (O4) is underway as of May 23, 2023111https://observing.docs.ligo.org/plan, and the search for gravitational wave events and their counterparts (Abbott et al., 2020a) has resumed. Searches for kilonovae are challenging due to the fact that they are short lived events, can be relatively faint, and may not be well localized. Sky localizations for gravitational wave events can span $\approx 100-10,000\ \mathrm{deg^{2}}$ (Röver et al., 2007; Fairhurst, 2009, 2011; Grover et al., 2014; Wen & Chen, 2010; Sidery et al., 2014; Singer et al., 2014; Berry et al., 2015; Essick et al., 2015; Cornish & Littenberg, 2015; Klimenko et al., 2016). However, in spite of these challenges, it is imperative to locate the transient as quickly as possible in the hope of observing the peak of emissions. A number of wide-field survey telescopes are used to try and cover these large sky localizations, such as: the Panoramic Survey Telescope and Rapid Response System (Pan-STARRS) (Morgan et al., 2012), Asteroid Terrestrial-impact Last Alert System (ATLAS) (Tonry et al., 2018), the Zwicky Transient Facility (ZTF) (Bellm et al., 2018; Graham et al., 2019; Ahumada et al., 2024), and in the near future BlackGEM (Bloemen et al., 2015), the Large Synoptic Survey Telescope (LSST) (Ivezić et al., 2019), the Nancy Grace Roman Space Telescope [^2] https://roman.gsfc.nasa.gov/ (Andreoni et al., 2024), and Ultraviolet Explorer (UVEX) (Kulkarni et al., 2021). While GW178017 is the only CBC merger event that has provided us with joint observations of a GW signal, kilonova, and short gamma-ray burst (sGRB), it is also potentially possible to identify kilonovae associated with sGRBs (Tanvir et al., 2013; Ascenzi et al., 2019; Jin et al., 2020; Rastinejad et al., 2022), or even serendipitously in survey operations (Andreoni et al., 2021; Almualla et al., 2021).
@@ 2.4 Dynamical and disk wind ejecta from fits: BNS and NSBH
$\dfrac{M^{dyn}_{\rm ej}}{10^{-3}M_{\odot}}=\mathrm{max}\left(0,\left[\dfrac{a}{C_{1}}+b+\left(\dfrac{m_{1}}{m_{2}}\right)^{n}+cC_{1}\right]m_{1}+[1\leftrightarrow 2]\right)$ (5)